OpenStudy (anonymous):

If f(x) = 5x + 1 then its inverse function is given by f^-1(x) = ...? Explainnn.

OpenStudy (anonymous):

\[y=f(x)\] \[y = 5x+1\] \[\frac{y-1}{5}= x\] Hence \[f^{-1}(x) = \frac{x-1}{5}\]

OpenStudy (angela210793):

x=5y+1 5y=x-1 y=(x-1)/5 f^-1=(x-1)/5

OpenStudy (anonymous):

To find \(f^{-1}(x)\) remember that \(f(f^{-1}(x)) = x\) Therefore: \[f(f^{-1}(x)) = 5(f^{-1}(x)) +1\]\[\implies x = 5f^{-1}(x) + 1\]\[\implies x-1 = 5(f^{-1}(x))\]\[\implies f^{-1}(x) = \frac{1}{5}(x-1)\]

OpenStudy (anonymous):

Or at least that's how I remember

OpenStudy (anonymous):

that makes absolutely no sence at all! I really don't understand f(x) at all. & when you add a ^-1 in there, it throws me off twice as much.